OpenStudy (anonymous):

Trevor solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 A)He should have substituted 5 + 2x B)He combined like terms incorrectly, it should have been 4x instead of 5x C)He subtracted 10 from the right side instead of adding 10 to the right side D)He made no mistake

OpenStudy (anonymous):

there is no way to make out the steps can you please write them again?

OpenStudy (anonymous):

sorry about that copying and pasting doesnt work to well

OpenStudy (mathstudent55):

2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 The mistake is in the next step. When you add 10 to both sides, you should get 5x = 20, but he got 5x = 0, which is wrong. x = 0 2(0) + y = 5 y = 5